OpenStudy (gliscorguy25):

Solve the absolute value equation. |x-3|=1. I need a refresher adn forgot how this works

OpenStudy (tkhunny):

Think on this: |1| = 1 |-1| = 1

OpenStudy (gliscorguy25):

yeah but does it matter that its supposed to be x-3 so would it be x+3?

OpenStudy (tkhunny):

Think on this: (x-3) = 1 -(x-3) = 1

OpenStudy (gliscorguy25):

alright thanks. sorry for so many questions

OpenStudy (tkhunny):

That's what we do is questions. (x-3) = 1 ==> x-3 = 1 ==> x = 4 -(x-3) = 1 ==> -x+3 = 1 ==> -x = -2 ==> x = 2 And we see the two solutions. Now, you do one!